# Explosive synchronization in networks of Type-I neurons with electrical synapses

### v1

@@ III.1 Networks of QIF neurons
wExploiting the mapping between the QIF neuronal network and the Kuramoto model Clusella, Pietras, and Montbrió (2022), we also investigated synchronization transition in partially degree-frequency correlated QIF networks, which are known to show ES in the Kuramoto networks Pinto and Saa (2015). To achieve partial degree-frequency correlation in the QIF network, a fraction of the neuron population is assigned frequencies following the $\eta_{i}=\bar{\eta}+\varepsilon k_{i}$ rule, while the rest are assigned random frequencies. Fig. 2 shows that the first-order transition to synchronization observed in a fully degree-frequency correlated SF network of QIF neurons persists even when only $10\%$ of the highest degree neurons are correlated. In contrast, when the frequencies of all the neurons are uncorrelated with their degrees, the transition becomes continuous.
Next, we describe transitions in star networks of fully degree-frequency correlated QIF neurons. We assign frequencies following $\eta_{i}=\bar{\eta}+\varepsilon k_{i}$ for every neuron. Simulations are performed by varying g following the same method as for the scale-free networks. Eq. (2) is solved using the RK4 method with step size dtdt ($\approx 10^{-4}$). Simulations are carried out till $T_{f}=25000$, and the order parameter is calculated over the final 10000 time units. Fig. 3 shows a sharp transition to synchronization across all network sizes. The forward transition points shift to lower values of $g$ as the size of the star network is increased.
Figure 3: Forward simulations of Star networks of QIF neurons with complete degree-frequency correlation. N is the number of neurons in the network. For these simulations $\bar{\eta}=20$, $\varepsilon=0.01$.
A network of QIF neurons with pure electrical coupling is known to reduce to the standard Kuramoto model Clusella, Pietras, and Montbrió (2022), yielding the following expression for its phase evolution.
$\dot{\theta_{i}}=\bigg(\frac{2\sqrt{\bar{\eta}}}{\tau}+\frac{\varepsilon k_{i}}{\tau\sqrt{\bar{\eta}}}\bigg)+\frac{\varepsilon g}{\tau}\sum_{j}^{N}\sin(\theta_{j}-\theta_{i})$ (7)
where $\omega_{i}=(\frac{2\sqrt{\bar{\eta}}}{\tau}+\frac{\varepsilon k_{i}}{\tau\sqrt{\bar{\eta}}})$ and $\lambda=\varepsilon g/\tau$ are the frequency and the coupling strength respectively of the reduced Kuramoto model. In Fig. 4, the results of simulating a complete degree-frequency correlated star network of 21 QIF neurons for different levels of heterogeneity characterized by $\varepsilon$ values are shown. In these simulations, when $g$ values are less than a certain value Leyva et al. (2025), a small randomness is added to the $\eta$ values during the backward run. This is done to slightly perturb the system out of the synchronized branch. These plots show that increasing $\varepsilon$ shifts the forward transition point ($\lambda_{cf}$) to a higher value. This trend aligns with the increase in the forward transition point with the rise in heterogeneity as reported for the Kuramoto network in W. Liu et al Liu et al. (2013). For a complete degree-frequency correlated star network with K leaves, the hub frequency is $\omega_{h}=K$, and the frequency of the leaves is $\omega_{l}=1$. The mean-field calculation carried out for such a network of Kuramoto oscillators gives the critical coupling strength $\lambda_{c}=(\Omega-\omega_{l})=\frac{K-1}{K+1}$ where $\Omega$ is the average frequency of the network Gómez-Gardeñes et al. (2011). Below this, a stable synchronized state does not exist. From the expression of $\omega_{i}$ in the Eq. (7), the average frequency of the reduced Kuramoto model will be $\Omega=(2\pi/\tau)+(2\varepsilon K/(\tau\bar{\eta}(K+1)))$. Thus, the corresponding theoretical $\lambda_{c}$ for the star network of QIF neurons is
For $\varepsilon=0.0095$ and $\varepsilon=0.0065$, theoretically predicted $\lambda_{c}$ values are $\approx 0.0019$ and $\approx 0.0013$ respectively. In Fig. (4), it is seen that the backward transitions occur at $\lambda$’s close to these values predicted by the above equation.
@@ III.2 Networks of Morris-Lecar neurons
In the previous section, we reported that the SF networks of the normal form of Type-I neurons (the QIF model) show explosive synchronization with complete and partial degree-frequency correlation. However, the transition becomes smooth if the degrees and the frequencies are uncorrelated. Motivated by these results, we consider Type-I Morris-Lecar neurons on the same network topologies with different levels of degree-frequency correlation. A Morris-Lecar neuron, modeled by Eq. (5), shows Type-I behavior for the chosen parameter values of section II. With these parameter values a SNIC bifurcation, separating excitable and oscillatory states, occurs at $I_{ext}\approx 39.6$ Tsumoto et al. (2006). Following section II, we correlate the frequency of a neuron with its degree as per the rule $I_{i,ext}=I_{0}+\varepsilon k_{i}$, with $I_{0}>39.6$. Simulations were performed for scale-free networks of 1000 ML neurons. The strength of the electrical synapse was varied following the procedure considered for scale-free networks of QIF neurons. The order parameter was calculated using the phases defined in Eq. (6).
@@
[…]
Figure 6: Plots of 1000 ML neurons connected through an SF network. In all cases, the same SF network with $\gamma=2.4$ is used. (a) Complete degree-frequency correlation. (b) Only the top $10\%$ of neurons with the highest degree are assigned frequencies correlated with their degree, while the rest are assigned random frequencies. (c) All neurons are assigned random frequencies. In both (b) and (c), random frequencies are drawn from a Lorentzian distribution $p(k)=a/\pi((k-k_{0})^{2}-a^{2})$ with $k_{0}=0$ and $a=1$. For all cases $I_{0}=43.0$ and $\varepsilon=0.01$.
Partial degree-frequency correlation is achieved by assigning frequencies to the uncorrelated neurons in the same way as was done for the QIF networks. Simulation results are shown in Fig. 6. We observe an abrupt transition to a synchronized state with only $10\%$ of correlation. When the neuron frequencies are completely uncorrelated to their degrees, the transition becomes continuous.
With these results of partial degree-frequency correlation simulations of ML neuronal network, we show that any Type-I neurons, with weak heterogeneity and coupling, can show ES if the neuronal network satisfies the conditions needed for ES in the corresponding Kuramoto model. These observations verify the mapping among these three models, which approximates the ML model with the QIF model, and the QIF model with the Kuramoto model.
@@ IV Conclusion
We have demonstrated the emergence of explosive synchronization in scale-free and star networks of Type-I neurons with electrical coupling and weak intrinsic heterogeneity. Our findings broaden the presence of ES, showing that it is not limited to specific neuronal models but can arise across a broad class of neurons. We establish this by demonstrating ES in the normal form of Type-I neurons (QIF model) and affirming it in a biophysical model of Type-I neurons (Morris-Lecar model). Our results in Fig. 1, 2 ,5 ,6 show that ES is robust to the degree exponent of scale-free networks and to various levels of degree-frequency correlations. For ES on star networks of QIF neurons, there was quantitative agreement between simulations and theory. Specifically, the critical coupling for backward transitions in numerical simulation agrees with theoretical predictions from the mapped Kuramoto model.
ES is hypothesized as the mechanism for epileptic seizures and the onset of anesthetic-induced unconsciousness. Understanding the conditions of explosive synchronization in neuronal networks could shed light on normal and pathological brain dynamics. Leveraging the mapping between Type-I neurons and the Kuramoto model, a large set of scenarios can be explored for explosive synchronization. This work is the first step towards it.

### v2

@@ III.1 Networks of QIF neurons
Fig. 5 shows that the first-order transition to synchronization observed in a fully degree-frequency correlated SF network of QIF neurons persists even when only $10\%$ of the highest degree neurons are correlated. In contrast, when the frequencies of all the neurons are uncorrelated with their degrees, the transition becomes continuous.
  We examine the effect of increasing the level of heterogeneity. As shown in Fig. 6, for a degree-frequency correlated SF network of QIF neurons hysteresis width progressively shrinks as $\varepsilon$ is increased and becomes negligible as $\varepsilon$ approaches 0.4. On further increasing $\varepsilon$, the sharpness in the transition decreases and tends to become continuous. In contrast, for the reduced QIF-Kuramoto model, the hysteresis width is unaffected by $\varepsilon$ (not shown). Thus, for sufficiently large heterogeneity, ES conditions for the Kuramoto model may not serve as a reliable guide for selecting ES conditions for the QIF model.
@@ III.2 Networks of Morris-Lecar neurons
In the previous section, we reported that the SF networks of the normal form of Type-I neurons (the QIF model) show explosive synchronization with complete and partial degree-frequency correlation. However, the transition becomes smooth if the degrees and the frequencies are uncorrelated. Now, we generalize these results to Type-I Morris-Lecar neurons operating close to the SNIC bifurcation on the same network topologies with different levels of degree-frequency correlation. A Morris-Lecar neuron, modeled by Eq. (II), shows Type-I behavior for the chosen parameter values of section II. With these parameter values an SNIC bifurcation, separating excitable and oscillatory states, occurs at $I_{ext}\approx 39.6\mu A/cm^{2}\$Tsumoto et al. (2006). Following section II, we correlate the frequency of a neuron with its degree as per the rule $I_{i,ext}=I_{0}+\varepsilon k_{i}$, with $I_{0}=43.0\mu A/cm^{2}$ and $\varepsilon=0.01$. With these parameters, $I_{ext}$ for nearly all neurons lies between $43.0\mu A/cm^{2}$ and $44.0\mu A/cm^{2}$ for scale-free networks considered in our simulations (see Appendix C for details), consistent with weak heterogeneity. Moreover, with this range of $I_{ext}$, all neurons lie close to the SNIC bifurcation. The Simulations were performed for scale-free networks of 1000 ML neurons. The strength of the electrical synapse was varied following the procedure considered for scale-free networks of QIF neurons. The order parameter was calculated using the phases defined in Eq. (7).
@@
Partial degree-frequency correlation is achieved by assigning frequencies to the uncorrelated neurons in the same way as was done for the QIF networks. Simulation results are shown in Fig. 8. We observe an abrupt transition to a synchronized state with only $10\%$ of correlation. When the neuron frequencies are completely uncorrelated to their degrees, the transition becomes continuous. With these results of ML neuronal networks, we show that any Type-I neurons, with weak heterogeneity and coupling, can show ES if the neuronal network satisfies the conditions needed for ES in the corresponding Kuramoto model.
These observations verify the mapping among these three models, which approximates the ML model with the QIF model, and the QIF model with the Kuramoto model.
@@ IV Conclusion
We have demonstrated the emergence of explosive synchronization in scale-free and star networks of Type-I neurons with weak electrical coupling and weak intrinsic heterogeneity. Specifically, we have shown the occurrence of ES in the normal form of Type-I neurons (QIF model) and generalized it in a biophysical model of Type-I neurons (Morris-Lecar model) close to the SNIC bifurcation. Our results in Fig. 4 and 7 show that ES is robust to the degree exponent of scale-free networks. Fig. 5 and 8 show that ES is observed for both complete and partial degree-frequency correlation, and the transition becomes continuous when there is no correlation. Thus, the explosive synchronization observed here is network-induced rather than a manifestation of the macroscopic dynamics observed through mean firing rate models.
The mapping between the QIF and the Kuramoto model Clusella, Pietras, and Montbrió (2022) served as a motivation for employing ES scenarios of the Kuramoto networks to our work. To validate this mapping between weakly heterogeneous and weakly coupled QIF and the Kuramoto model, we compared the simulation and analytical results of the reduced QIF-Kuramoto model with the corresponding degree-frequency correlated star network of QIF neurons as shown in Fig. 1 and 2. The quantitative agreement was close, specifically for the critical coupling of backward transitions. At strong heterogeneity, the mapping does not appear to hold as seen in our simulations of SF networks of QIF neurons with increasing heterogeneity. As shown in Fig. 6, hysteresis vanishes, and the transition tends to become continuous as heterogeneity increases.
Our findings broaden the presence of ES, showing that it is not limited to specific neuronal models but can arise across a broad class of neurons. ES is hypothesized as the mechanism for epileptic seizures and the onset of anesthetic-induced unconsciousness. Understanding the conditions of explosive synchronization in neuronal networks could shed light on normal and pathological brain dynamics. Leveraging the mapping between Type-I neurons and the Kuramoto model, a large set of scenarios can be explored for explosive synchronization. This work is the first step towards it.
@@
Appendix A Choosing Threshold voltages ($V_{p}$) for the QIF model
In our simulations, phases of QIF neurons are computed using Eq. (3), while Eq. (4) employs phases computed in the asymptotic limit of $V_{p}\rightarrow\infty$. In this appendix, we show that the two phases approach each other as the threshold voltage is increased in Eq. (3). We select a $V_{p}$ and consider 5000 uniformly distributed voltages between $V_{p}$ and $V_{r}$ (with $V_{r}=-V_{p}$). At these voltage values, we calculate the two phases and compute the mean and absolute differences. Fig. 9 shows that as the threshold voltage is increased, the phase of Eq. (3) tend towards the asymptotic phase indicated by the decreasing mean and the maximum error. This serves as a reference for selecting $V_{p}=750$ in our numerical simulations.
Figure 9: Increasing $V_{p}$ values decreases the error between the phase in Eq. (3) and the asymptotic phase. Here $\eta=20$ is taken in Eq. (3)
Appendix B Analytical expression for the critical coupling of degree-frequency correlated star network of QIF neurons
The mean-field analysis of the Kuramoto model on a star network with K leaves results in a critical coupling
$\lambda_{c}=(\Omega-\omega_{l})$ (9)
where the synchronized state loses stability Gómez-Gardeñes et al. (2011). Here, $\Omega$ and $\omega_{l}$ are the average frequency of the network and the common frequency of the leaves, respectively. Eq. (9) is obtained by considering phases in a frame rotating with frequency $\Omega$ and imposing the phase-lock conditions on the hubs and leaves. Furthermore, for a degree-frequency correlated star network with $\omega_{l}=1$ and $\omega_{h}=K$ as the frequencies of the leaves and hubs, respectively, it has been shown that Gómez-Gardeñes et al. (2011) Eq. (9) reduces to
$\lambda_{c}=\frac{K-1}{K+1}$ (10)
For the degree-frequency correlation considered in our work in the reduced QIF-Kuramoto model, Eq. (4), frequencies of the hub and leaves are {subequations} {align} ω_h =2¯ητ + εKτ¯η
ω_l =2¯ητ + ετ¯η The resulting average frequency is
$\Omega=(2\sqrt{\bar{\eta}}/\tau)+(2\varepsilon K/(\tau\bar{\eta}(K+1)))$ (11)
Substituting the frequencies Eq. (11) and Eq. (10) into Eq. (9) gives the critical coupling of the reduced QIF-Kuramoto model as
This is the theoretical $\lambda_{c}$ for the star network of the reduced QIF-Kuramoto model, as mentioned in Eq. (8).
Figure 10: Distribution of external current of degree-frequency correlated Morris-Lecar neurons with scale-free network ($\gamma=2.4$) considered in Fig. 8. $99\%$ of the neuron fraction has $I_{ext}$ between 43.0 and 44.0
Appendix C $I_{ext,i}$ distribution in Morris-Lecar scale-free neuronal network
We consider the scale-free network with $\gamma=2.4$ (used in Fig. 8) and plot the distribution of external current $I_{ext,i}$ when the neurons are fully degree-frequency correlated following the $I_{ext,i}=I_{0}+\varepsilon k_{i}$ rule. Fig. 10 shows that the external current distribution is peaked very close to the $I_{0}$ value, and $99\%$ of the neuron population has $43.0\mu A/cm^{2}<I_{ext}<44.0\mu A/cm^{2}$. Therefore, in our simulations, neurons’ intrinsic dynamics are weakly heterogeneous.